Two small spheres spaced apart have equal charge. How many electrons must be present on each sphere if the magnitude of the force of repulsion between them is 4.57 x 10^-21 N?


1
Expert's answer
2022-02-17T11:37:07-0500

Answer

No. Of electrons must be present on each sphere

"n=\sqrt{\frac{Fr^2}{ke^2}}"

Take value of separation between them is r=1m

"n=\sqrt{\frac{4.57*10^{-21}*1^2}{9*10^9*1.6*1.6*10^{-38}}}\\n=4450"